answersLogoWhite

0

🎒

Probability

The probability of a certain event is a number expressing the likelihood that a specific event will occur, expressed as the ratio of the number of actual occurrences to the number of possible occurrences. In mathematics, it is a measure of how often an event will happen and is the basis of statistics.

14,643 Questions

What is the probability of getting tails twice in a row?

The probability of tossing a coin twice and getting tails both times is 1 in 4, or 25%.

If you have already tossed a coin and had it land on tails, the probability that it will land on tails again the next time you toss it is 50%.

What does square of probability explain?

If p is the probability that an event will happen once, then the probability that it will happen just twice is p2. The probability it will happen 3 times is p3. The probability it will happen at least once ( ie once or twice or three times ore more times is p + p2 + p3 + ... = p(1-p). For "or" you add probabilities, for "and" you multiply probabilities.

How do you compute C11 5 the 11 and 5 are subscripts?

I'm assuming this is a combinations problem.

nCk = n!/(k!(n-k)!) is the general equation.

In your case n= 11 and k = 5

for 11!/(5!(11-5)!) = 11!/(5!*6!) = (11*10*9*8*7)/(5*4*3*2*1) = 462

What is the probability of being dealt a spade and heart in a deck of 52 cards?

Number of cards in a deck = 52

Number of cards that are spade = 13

Number of cards that are heart = 13

Probability that the card drawn is a spade and heart =

13/52 + 13/52

= 26/52 or 1/2

What does the square root of 3 equal to?

1.73205081The question could be restated, what number times itself equals 3. To the 8th decimal, the answer is 1.73205081 x 1.73205081 = 3

Why can't 120 percent be a probability of an event?

The probability of an even must be a real number in the range [0, 1]. 120% is outside that range.

What does equally unlikely mean in probability?

Two events are equally unlikely if the probability that they do not happen is the same for each event. And, since the probability of an event happening and not happening must add to 1, equally unlikely events are also equally likely,

What best describes what the mishap probability component is?

the mishap probability component is an estimation that a hazard will result in a mishap

How many simple random samples of size 5 can be selected from a population of size 34?

278 256

The number of 5 different item combinations from a pool of 34 different items is

given by:

34C5 = 34!/(5!29!) = 278 256

When two dice are rolled what is the probability of getting a sum of 5 or 6?

When two dice are rolled, the probability of getting:

2 is 1/36

3 is 2/36

4 is 3/36

5 is 4/36

6 is 5/36

7 is 6/36

8 is 5/36

9 is 4/36

10 is 3/36

11 is 2/36

12 is 1/36

Why is the standard error a smaller numerical value compared to the standard deviation?

Let sigma = standard deviation.

Standard error (of the sample mean) = sigma / square root of (n), where n is the sample size. Since you are dividing the standard deviation by a positive number greater than 1, the standard error is always smaller than the standard deviation.

What does it mean when the standard error value is smaller than the standard deviation?

It simply means that you have a sample with a smaller variation than the population itself. In the case of random sample, it is possible.

What is the probability that you will get a 3-digit combination correct if no digits in the combination can be repeated?

If the digit 0 is allowed then the answer is 1/(10*9*8) = 1/720 = 0.0014 approx.

If 0 is disallowed, then 1/(9*8*7) = 1/504 = 0.0020 approx.

What is the probability of obtaining 2 ones in a row when rolling a fair six sided die?

If you roll the die often enough, the probability of getting a 2 ones in a row is 1.

In only 2 rolls, the answer is (1/6)*(1/6) = 1/36

What is the parameters that determine a Binomial Distribution?

The binomial distribution has two parameter, denoted by n and p.

n is the number of trials.

p is the constant probability of "success" at each trial.

What is the probability of drawing a heart a diamond or a spade?

the probibility of drawing a heart diamond or a spaid is a 1/3 chance........

What is the probability of drawing an eight or a spade out of a deck of 52 cards?

Number of cards in a deck =52

Number of cards that are spade =13

Number of cards that are eight =4

Number of cards that are spade and eight =1

Probability that the card drawn is a spade or an eight =

13/52+4/52 -1/52

=16/52 or 4/13